OpenStudy (anonymous):

Simplify: (1+cos2x)/sin2x

OpenStudy (primeralph):

is it cos^2(x) or cos(2x)

OpenStudy (anonymous):

@elly394 Can you answer primer's question?

OpenStudy (anonymous):

it is cos(2x)

OpenStudy (jhannybean):

@FutureMathProfessor , @bahrom7893

OpenStudy (bahrom7893):

What is Cos(2a) ?

OpenStudy (bahrom7893):

Okay, I guess I'll answer it: Cos(2a) = Cos(a+a) = Cos(a)Cos(a) - Sin(a)Sin(a) = Cos^2(a) - Sin^2(a)

OpenStudy (jhannybean):

Ohh!! nevermind. i did cos^(2) x -sin^(2) x

OpenStudy (bahrom7893):

We also know that 1 = Sin^2(a) + Cos^2(a), so top becomes: Sin^2(a) + Cos^2(a) + Cos^2(a) - Sin^2(a) = 2Cos^2(a)

OpenStudy (jhannybean):

HAHAHA spoke too soon.

OpenStudy (bahrom7893):

I mean it doesn't matter if you have x or a. I wonder how my xs somehow turned into as.

OpenStudy (bahrom7893):

And on the bottom we have: Sin(2a) = 2Sina*Cosa

OpenStudy (jhannybean):

they didn't :P

OpenStudy (jhannybean):

ohh...soyou can cancel out a cos(a)

OpenStudy (bahrom7893):

2Cos^2(a)/(2Sina*Cosa) = Cosa/Sina = Tanx

OpenStudy (bahrom7893):

I mean cotangent x

OpenStudy (bahrom7893):

and somewhere along the way my xs became as and then xs again.

OpenStudy (jhannybean):

yesss :D (2cos^(2) (a)) / (2 sin(a) * cos(a)) = cos (a)/sin(a) = cot(a)